OpenStudy (kimo_indie):

Which value of y forms the solution of the system defined by y = 11x + 1 and 11x + 12y = 12? a) 1 b) 1/13 c) 11 d) 11/13

OpenStudy (jahmiya):

y = 11x+1 11x+12y=12 is the same as y = -(11x/12) + 1 y=11x+1 y=(-11x/12)+1 multiply both sides by 12 to get rid of the fraction 12y=-11x+12 y = 11x+1 12y = -11x +12 add them together: 13y = 0x +13 13y = 13 divide both sides by 13 y = 1